# Identifying compact symmetric objects with high-precision VLBI and Gaia astrometry

### v1

@@ Abstract
Context. Compact symmetric objects (CSOs) represent a key early stage in the evolution of radio galaxies, but their reliable identification is challenging due to the difficulty of detecting their radio cores and distinguishing them from other compact radio sources.
Aims. We develop and validate a new method to unambiguously identify CSOs by combining Gaia’s high–precision optical astrometry with the high-resolution radio imaging of very long baseline interferometry (VLBI).
Methods. We analyze 40 CSO candidates by overlaying Gaia DR3 positions on VLBI maps to pinpoint their central engines. CSOs are confirmed when their Gaia positions lie between symmetric radio lobes. In contrast, core–jet sources are identified when the optical position coincides with one end of the radio structure. To verify CSO classifications, we also analyze spectral index maps, source variability, and jet kinematics using multi–epoch VLBI observations spanning up to 25 years.
Results. Our method successfully identified 22 genuine CSOs and 10 core–jet sources, while 8 other objects remain ambiguous due to significant optical–radio positional offsets. The confirmed CSOs display diverse properties, with kinematic ages ranging from 20 years to over 1000 years, and hotspot advance speeds typically below , although some sources show mildly relativistic expansion. Their spectral and variability characteristics support classification as young, evolving radio galaxies. Notably, five nearby CSOs show systematic optical–radio offsets despite strong CSO morphological evidence, highlighting the influence of the host galaxy environment.
Conclusions. The Gaia-VLBI method provides a powerful new tool for reliably identifying CSOs. Our sample reveals a diverse distribution of radio powers, confirming both high-power and low-power CSOs and suggesting the existence of multiple evolutionary paths. High-power CSOs may evolve into large-scale radio galaxies, while many low-power CSOs show evidence of being confined by their host galaxy environments. Our findings suggest that CSO evolution is not only affected by the intrinsic jet power but also by environment, the duration and sustainability of the central engine activity. Future high-sensitivity observations of low-power CSOs are crucial for understanding the full range of formation mechanisms and evolutionary paths of radio galaxies.
@@
Compact symmetric objects (CSOs) represent a key early stage in radio galaxy evolution, characterized by symmetric double-lobed morphology within a projected size of kiloparsec (kpc) (Phillips & Mutel, 1980, 1982; Pearson & Readhead, 1981, 1988). These compact radio sources offer unique insights into the complex interplay between active galactic nuclei (AGN) activity and their host environments and have become laboratories for studying jet–ISM interactions and AGN feedback mechanisms (Tadhunter, 2016; Bicknell et al., 2018; O’Dea & Saikia, 2021).
The physical nature of CSOs has been debated through three primary theoretical frameworks. The Young Radio Galaxy Model suggests CSOs represent the earliest evolutionary stage of radio galaxies (e.g. Phillips & Mutel, 1980; Readhead et al., 1996a; Owsianik & Conway, 1998), supported by the observed “luminosity–size” relationship (Baldwin, 1982) that indicates progression from CSOs to larger-scale radio sources (e.g. Fanti et al., 1995; Snellen et al., 2000; Kunert-Bajraszewska et al., 2010; An & Baan, 2012). However, this evolutionary path may account for only a small fraction of the overall CSO population, primarily those with high-power jets, while the fate of those with lower-power and unstable jets is uncertain.
The Frustrated Radio Galaxy Model proposes that CSOs are impeded by dense interstellar medium (ISM) in their host galaxies (e.g. van Breugel et al., 1984; Wilkinson et al., 1984; O’Dea et al., 1991; De Young, 1993), particularly relevant for low-radio-power sources that may never evolve into Fanaroff–Riley type I or II radio galaxies (Fanaroff & Riley, 1974). Recent observations of FR 0 radio galaxies (Baldi et al., 2018, 2019) and even lower-power radio sources (Chilufya et al., 2024) suggest a potential link to previously unexplored low-power CSO population, where jet–ISM interactions significantly shape radio morphology.
@@
The Transient Phenomenon Model suggests some CSOs may be short-lived, resulting from intermittent or episodic AGN activity. Here, it is important to distinguish between “intermittent” activity (repeating cycles caused by accretion disk instabilities, e.g., Janiuk et al. 2002; Czerny et al. 2009; Siemiginowska et al. 2010; Wu 2009) and “episodic” events (one-time phenomena such as tidal disruption events, TDEs, Readhead et al. 1996b, 2024). While this episodic TDE model can explain certain short-lived sources, it struggles to account for CSOs that evolve into larger radio galaxies or show multiple activity episodes (e.g. Schoenmakers et al., 2000; Kaiser et al., 2000; Marecki & Szablewski, 2009; Joshi et al., 2011; Mahatma et al., 2019; Orienti et al., 2023). The presence of double-double radio lobes or remnant emission suggests that CSOs may undergo recurrent activity cycles (Machalski et al., 2010; Konar & Hardcastle, 2013), necessitating models that can explain both transient and long-term evolutionary processes.
@@
To connect our understanding of radio galaxy formation, precise quantification of CSO dynamics, ages, and triggering rates is crucial to establish their relationship to large-scale radio galaxies. CSOs provide unique insights into the birth and evolution of radio galaxies through several key aspects: their compact size and youth provide direct windows into early-stage jet activity and jet–ISM interactions, while their symmetric double-lobed morphology, typically oriented close to the sky plane, minimizes relativistic beaming effects and reveals intrinsic jet properties.
@@
CSOs are particularly critical to understanding AGN feedback mechanisms and galaxy evolution. They serve as laboratories for studying three critical processes: initial jet–ISM coupling in galactic nuclei, jet physics with minimal projection effects, and diverse evolutionary pathways spanning from high-power sources evolving into large radio galaxies to confined low-power sources. These characteristics make CSOs fundamental to understanding how AGN feedback operates in practice and influences galaxy evolution across cosmic time.
@@
Despite their importance, the identification and physical nature of CSOs remain challenging areas of investigation. The primary difficulty lies in detecting their radio cores, which are often weak or obscured (Gugliucci et al., 2005) due to several factors. First, minimal relativistic beaming effects due to large inclination angles result in reduced core emission (Peck et al., 1999; Orienti et al., 2006; Marr et al., 2001; Orienti & Dallacasa, 2008; An et al., 2012; Wu et al., 2013). Second, dense ionized gas in nuclear regions can cause significant attenuation through synchrotron self-absorption or free-free absorption (O’Dea, 1998; Fanti, 2009). Third, frequency-dependent opacity effects can shift apparent core positions, complicating their identification in VLBI observations.
@@
Traditional CSO identification has relied heavily on morphological and spectral criteria derived from high-resolution very long baseline interferometry (VLBI) observations. Key indicators include symmetric double-lobed structure, steep-spectrum lobes, and low variability (e.g., Taylor et al., 1996; Peck & Taylor, 2000; Healey et al., 2007). However distinguishing genuine CSOs from core–jet sources exhibiting similar structures remains difficult in the absence of clear core detection. While radio spectral indices from multi-frequency VLBI observations provide supplementary evidence, uncertainties in the CSO variability, dynamics and life cycles limit our ability to construct complete, unbiased CSO samples.
@@
To address these challenges, we introduce a new method that combines high-precision optical astrometry of the Gaia space mission (Gaia Collaboration et al., 2016) with high-resolution VLBI imaging to unambiguously identify genuine CSOs. Gaia Data Release 3 (DR3, Gaia Collaboration et al., 2023) provides unprecedented astrometric precision (0.04 mas for bright sources with apparent optical magnitude , and mas at ) for approximately 1.6 million quasar-like objects. When overlaid on VLBI maps, these precise optical positions can effectively pinpoint AGN central engines, even when radio cores are weak or undetectable.
@@
This study applies this method to 40 CSO candidates selected from existing VLBI surveys (see details in Section 2). By overlaying Gaia DR3 positions on multi-epoch, multi-frequency VLBI maps, we can definitively classify these objects and study their properties in unprecedented detail.
@@
The structure of this paper is as follows: Section 3 presents the identification of CSOs and core–jet sources using the so-called Gaia+VLBI method. Section 4 discusses the radio properties of the confirmed CSOs. The results are summarized in Section 5. Throughout this paper, we assume a standard flat Cold Dark Matter cosmology with km s-1 Mpc-1, , and .
@@ 2.1 CSO identification criteria
Traditional identification of CSOs relies on several key characteristics that distinguish them from other compact radio sources. These diagnostic criteria include:
Compact and symmetric radio structure: projected linear size kpc; double lobes with symmetric brightness distribution (brightness ratio 10:1); lobes positioned on opposite sides of a central core, forming a mirror-symmetric appearance.
Spectral properties: sources typically exhibit convex or steep radio spectrum, peaking at GHz frequencies; lobes have steep spectra (with power-law spectral index , where ); the core has a flat or inverted spectrum.
Low variability: typical variability level ;
Sub-relativistic hotspot advance speeds: generally , where denotes the speed of light.
However, comprehensive data satisfying all these criteria simultaneously remains challenging to obtain for most candidate sources, necessitating new approaches to CSO identification.
@@ 2.2 VLBI CSO sample
The COINS (CSOs Observed in the Northern Sky) sample contributed 52 CSOs and candidates from a VLBI continuum survey of approximately 2000 compact radio sources (Peck & Taylor, 2000). An & Baan (2012) provided 46 CSOs and candidates with kinematic measurements for 24 sources, enabling statistical analysis of their dynamical properties. The VLBA Imaging and Polarimetry Survey (VIPS) at 5 GHz yielded 103 CSO candidates from observations of 1127 sources (Taylor et al., 2005; Helmboldt et al., 2007), from which Tremblay et al. (2016) identified 24 CSOs and 33 new candidates.
Our CSO candidate sample combines data from these three CSO studies, yielding over 140 candidates based on symmetric morphology and steep or convex radio spectra (Appendix A). After removing redundant sources (Fig. 1), we established a consolidated sample of 105 CSO candidates (Appendix B). While not comprehensive in sky coverage or flux density limits, this dataset comprises the most credible CSOs and candidates confirmed through multiple criteria (primarily morphology and radio spectra) in previous studies, providing an effective testbed for our CSO identification technique.
@@ 2.3 Gaia+VLBI sample
We introduce a novel approach that combines high-precision optical astrometry from Gaia with high–resolution VLBI imaging to unambiguously identify CSO cores. While previous studies comparing Gaia and VLBI astrometry focused mainly on radio–-optical offsets in core–-jet sources (Petrov & Kovalev, 2017), our work represents the first systematic applications of this methodology to a large CSO sample, directly addressing the persistent challenge of identifying CSO cores.
By cross-matching our VLBI sample with the Gaia DR3 catalogue using a search radius, we identified 40 optical counterparts (Table 1) from 105 CSO candidates (Appendix C). The Gaia optical centroid corresponds to the immediate vicinity of the accretion disk and central supermassive black hole (SMBH), providing a reliable reference point for the AGN core position. When combined with VLBI’s sub-milliarcsecond (mas) resolution, this allows precise localization of the central engine in these young radio sources, even when the radio core is faint or obscured. The modest match rate (38%) suggests that most CSO nuclei are optically faint or obscured.
The cross-matched sample includes a heterogeneous mix of quasars (), radio galaxies, and other AGN types (Table 1 and Fig. 3). Most CSOs and candidates cluster at low and moderate redshifts (), while core–jet sources show a broader redshift range. This methodological approach offers significant advantages over traditional techniques:
precise core localization even when radio emission is weak or obscured
unambiguous discrimination between CSOs and core–jet sources
independent verification of radio structure orientation
enhanceed accuracy in kinematic measurements by establishing a fixed reference point
The combination of Gaia’s unprecedented astrometric precision ( mas for bright sources, mas at ) with VLBI’s sub-mas resolution creates a powerful tool for studying the birth and evolution of radio galaxies.

### v2

@@ Abstract
Context. Compact symmetric objects (CSOs) trace the earliest phases of radio-galaxy growth; however, robust classification is difficult when radio cores are weak or invisible.
Aims. We aim to develop and test a Gaia+VLBI approach that utilizes the high-precision optical astrometry of Gaia together with the high-resolution imaging of very long baseline interferometry (VLBI) to reliably locate the central engine and classify CSOs.
Methods. We analysed 40 literature CSO candidates by overlaying Gaia Data Release 3 (DR3) positions on VLBI maps and by examining spectral index distributions, whole-source variability, and hotspot kinematics over up to 25 years. A source is classified as a CSO when the Gaia centroid lies between two steep-spectrum lobes; alignment with one end of the radio structure indicates a core–jet source.
Results. Our method yields 20 confirmed CSOs, ten core–jet sources, and ten ambiguous cases affected by significant optical–radio positional offsets or limited data. The confirmed CSOs show low integrated variability, hotspot advance speeds typically , where denotes the speed of light (with a few mildly relativistic cases), and kinematic ages of yr. Five nearby CSOs show systematic Gaia–VLBI offsets despite the CSO-like morphology, likely reflecting host-galaxy environments and Gaia astrometric systematics. We find a clear dichotomy with radio power: high-power CSOs () tend to be larger and host faster hotspots, while many low-power systems remain sub‑kiloparsec and environmentally confined.
Conclusions. Gaia+VLBI registration is a powerful method for CSO classification, especially where radio cores are faint. The observed power-size-velocity-age relations support distinct multiple evolutionary tracks, with high-power CSOs plausibly growing into large radio galaxies, while low-power CSOs appear confined by their host galaxy environments. Taken together, our results indicate that CSO evolution is shaped not only by intrinsic jet power, but also by host–galaxy environment and the duty cycle of the central engine. High-sensitivity observations of low-power CSOs will be crucial to map the full diversity of formation channels and evolutionary pathways of radio galaxies.
@@
Compact symmetric objects (CSOs) represent a key early stage in radio galaxy growth, characterized by mirror-symmetric double-lobed morphology within a projected size of kiloparsec (kpc) (Phillips & Mutel, 1980, 1982; Pearson & Readhead, 1981, 1988). These sub-kiloparsec, symmetric radio sources provide clean laboratories for jet–interstellar medium (ISM) coupling and nascent active galactic nucleus (AGN) feedback with minimal projection effects (Tadhunter, 2016; Bicknell et al., 2018; O’Dea & Saikia, 2021).
@@
The physical nature of CSOs has been framed by three primary scenarios, which are as follows. (I) The young radio galaxy scenario suggests CSOs represent the earliest evolutionary stage of radio galaxies (e.g., Phillips & Mutel, 1980; Readhead et al., 1996a; Owsianik & Conway, 1998), supported by the observed “luminosity–size” relationship (Baldwin, 1982), which indicates progression from CSOs to larger scale radio sources (e.g., Fanti et al., 1995; Snellen et al., 2000; Kunert-Bajraszewska et al., 2010; An & Baan, 2012). However, this evolutionary path may only account for a small fraction of the overall CSO population, primarily those with high-power jets (), while the fate of those with lower power and unstable jets is uncertain. (II) In the frustration model, CSOs are confined by dense ISM in their host galaxies (e.g., van Breugel et al., 1984; Wilkinson et al., 1984; O’Dea et al., 1991; De Young, 1993), especially at low radio power, potentially linking them to Fanaroff–Riley (FR) type 0 (FR0) radio galaxies (Fanaroff & Riley, 1974; Baldi et al., 2018, 2019). Even lower power radio sources (Chilufya et al., 2024) suggest a potential link to a previously unexplored low-power CSO population, where jet–ISM interactions significantly shape radio morphology. (III) The transient-phenomenon model suggests some CSOs may be short-lived, resulting from intermittent or episodic AGN activity. It is important to distinguish between “intermittent” (e.g., repeating cycles caused by accretion-disk instabilities, Janiuk et al. 2002; Czerny et al. 2009; Siemiginowska et al. 2010; Wu 2009) and “episodic” events (one-time phenomena such as tidal disruption events, TDEs, Readhead et al. 1996b, 2024). While the episodic TDE model can explain certain short-lived sources, it struggles to account for CSOs that evolve into larger radio galaxies or show multiple activity episodes (e.g., Schoenmakers et al., 2000; Kaiser et al., 2000; Marecki & Szablewski, 2009; Joshi et al., 2011; Mahatma et al., 2019; Orienti et al., 2023). The presence of double-double radio lobes or remnant emission suggests that CSOs may undergo recurrent activity cycles (Machalski et al., 2010; Konar & Hardcastle, 2013), requiring models that explain both transient and long-term evolutionary processes.
@@
To understand radio-galaxy formation, precise quantification of CSO dynamics, ages, and triggering rates is crucial to establish their relationship with large-scale radio galaxies. CSOs provide unique insights into the birth and growth of radio galaxies through several key aspects: their compact size and youth provide direct windows into early-stage jet activity and jet–ISM interactions, while their symmetric double-lobed morphology, oriented close to the sky plane, minimizes relativistic beaming effects and reveals intrinsic jet properties.
@@
Compact symmetric objects are essential for understanding AGN feedback mechanisms and galaxy evolution. They serve as laboratories for studying three critical processes: initial jet–ISM coupling in galactic nuclei, jet physics with minimal projection effects, and diverse evolutionary pathways spanning from high-power sources evolving into large radio galaxies to confined low-power sources. These characteristics make CSOs crucial to understanding how AGN feedback operates and influences galaxy evolution across cosmic time.
@@
Despite their importance, identifying and understanding the physical nature of CSOs remains challenging. The primary difficulty lies in detecting their often weak or obscured radio cores (Gugliucci et al., 2005), caused by several factors. First, minimal relativistic beaming due to large inclination angles reduces core emission (Peck et al., 1999; Orienti et al., 2006; Marr et al., 2001; Orienti & Dallacasa, 2008; An et al., 2012; Wu et al., 2013). Second, dense ionized gas in nuclear regions can cause significant attenuation through synchrotron self-absorption or free-free absorption (O’Dea, 1998; Fanti, 2009). Third, frequency-dependent opacity effects can shift apparent core positions, complicating identification.
@@
Traditional CSO identification has relied on morphological and spectral criteria derived from high-resolution, very-long-baseline-interferometry (VLBI) observations. Key indicators include a symmetric double-lobed structure, steep-spectrum lobes, and low variability (e.g., Taylor et al., 1996a; Peck & Taylor, 2000; Healey et al., 2007). However, distinguishing genuine CSOs from core–jet sources exhibiting similar structures remains difficult in the absence of clear core detection. While multifrequency VLBI observations provide supplementary evidence, uncertainties in CSO variability, dynamics and life cycles limit our ability to construct unbiased CSO samples.
@@
To address these challenges, we introduce a new method combining high-precision optical astrometry from the Gaia space mission (Gaia Collaboration et al., 2016) with high-resolution VLBI imaging to reliably identify genuine CSOs. Gaia Data Release 3 (DR3, Gaia Collaboration et al., 2023) provides unprecedented astrometric precision (0.04 mas for bright sources with apparent optical magnitude and mas at ) for approximately 1.6 million quasar-like objects. When overlaid on VLBI maps, these precise optical positions pinpoint AGN central engines, even when radio cores are weak or undetectable.
@@
This study applies this method to 40 CSO candidates selected from existing VLBI surveys (Section 2). By overlaying Gaia DR3 positions on multi-epoch, multifrequency VLBI maps, we were able to classify these objects and study their properties in detail.
@@
The structure of this paper is as follows. Section 2 presents the methodology and parent sample. Section 3 presents the identification of CSOs and core–jet sources using the Gaia+VLBI method. Section 4 discusses the radio properties of the confirmed CSOs. The results are summarised in Section 5. Throughout this paper, we assume a standard flat cold dark matter cosmology with km s-1 Mpc-1, , and . The radio spectrum is defined as , where is the flux density at the observing frequency, ; and is the spectral index.
@@ 2.1 CSO identification criteria
We used four diagnostics for CSO identification:
1.
Morphology. A compact, mirror-symmetric, double-lobed radio structure with projected linear size of kpc and a lobe-brightness ratio of 10:1;
2.
Spectral properties. Steep-spectrum lobes and a flat or inverted-spectrum core (when detected); convex or overall steep radio spectra are common;
3.
Low integrated variability. Whole-source variability is typically ;
4.
Hotspot kinematics. Sub-relativistic hotspot advance speeds, typically , where denotes the speed of light.
In principle, a confirmed CSO satisfies all four criteria. In practice, items (3) and (4) require homogeneous, long-term monitoring and are not always available. We therefore used (1) and (2) for the initial classification and treated (3) and (4) as confirming tests where data exist. If any diagnostic conflicted with CSO expectations, we adopted the conservative label of CSO candidate (CSOc). Table 1 summarises, for each source, the diagnostics available and their outcomes, including the whole-source variability measurement.
| Name | Cl. | Gaia+VLBIaaaa = Gaia centroid near the midpoint between two lobes; = “large optical–radio offset.” | M.+S.bbbb is applied when a two-sided, edge-brightened structure with steep-spectrum lobes and flat or inverted core is established in VLBI maps. | ccccWhole-source variability at 2.3 GHz using matched () coverage; “” if . | dddd“” if hotspot separation speed is ; “” if ; “/” if not measured. |
| (1) | (2) | (3) | (4) | (5) | (6) |
| J0048+3157 | CSO | / | / | ||
| J0741+2706 | CSO | / | / | ||
| J0832+1832 | CSO | 0.044 () | |||
| J1110+4817 | CSO | 0.042 () | eeeeValues are limits or pattern-/orientation-affected; see kinematics discussion in Section 4. | ||
| J1111+1955 | CSO | 0.036 () | |||
| J1158+2450 | CSO | 0.046 () | |||
| J1234+4753 | CSO | 0.036 () | |||
| J1244+4048 | CSO | 0.024 () | eeeeValues are limits or pattern-/orientation-affected; see kinematics discussion in Section 4. | ||
| J1247+6723 | CSO | 0.032 () | |||
| J1256+5652 | CSO | 0.062 () | / | ||
| J1310+3403 | CSO | / | / | ||
| J1326+3154 | CSO | / | / | ||
| J1358+4737 | CSO | 0.002 () | |||
| J1407+2827 | CSO | 0.179 () | |||
| J1511+0518 | CSO | 0.077 () | |||
| J1602+2418 | CSO | 0.085 () | eeeeValues are limits or pattern-/orientation-affected; see kinematics discussion in Section 4. | ||
| J1755+6236 | CSO | / | / | ||
| J1815+6127 | CSO | 0.093 () | /eeeeValues are limits or pattern-/orientation-affected; see kinematics discussion in Section 4. | ||
| J1945+7055 | CSO | 0.109 () | |||
| J2355+4950 | CSO | / | / | ||
| J0119+3210 | CSOc | / | ggggfootnotemark: | ||
| J0650+6001 | CSOc | / | |||
| J0831+4608 | CSOc | / | / | ||
| J0906+4636 | CSOc | / | / | ||
| J0943+1702 | CSOc | / | |||
| J1148+5924 | CSOc | / | / | ||
| J1254+1856 | CSOc | / | / | ||
| J1309+4047 | CSOc | / | / | ||
| J1559+5924 | CSOc | / | / | ||
| J1823+7938 | CSOc | / |
@@ 2.2 VLBI CSO sample
Our CSO candidate sample combines data from three representative CSO studies: the COINS (CSOs observed in the northern sky) sample Peck & Taylor (2000), the A&B2012 sample An & Baan (2012), and the T2016 sample Tremblay et al. (2016). These studies yielded over 140 candidates based on symmetric morphology and steep or convex radio spectra. After removing redundant sources (Fig. 1), we established a consolidated sample of 105 CSO candidates. While not comprehensive in sky coverage or flux-density limits, this dataset represents the most credible collection of CSOs and candidates confirmed through multiple criteria (primarily morphology and radio spectra), providing an effective testbed for our CSO identification technique. The sample spans a broad range of redshifts (0.015 to 1.791, median 0.269; Fig. S1 in Supplementary Material222Additional figures, tables, and methods are available at Zenodo (https://doi.org/10.5281/zenodo.17278723). Figure and table numbering in Supplementary Material follows the S-scheme (e.g., Fig. S1, Table S1)).
We first selected 40 CSO candidates using Gaia+VLBI registration (Fig. 1), restricted to sources that already satisfy the morphological and radio spectral criteria. For objects with variability and/or hotspot motion measurements, any robust inconsistency with these diagnostics leads to a conservative downgrade to a CSOc. Some sources show large Gaia–VLBI offsets; however, strong independent-literature evidence supports their classification as CSOs (Table 1). Conversely, some objects show good Gaia–VLBI alignment, yet fail the low-variability or sub-relativistic hotspot speed expectations for standard CSOs, and are therefore tentatively listed as CSOc. We note minor discrepancies with classifications in Readhead et al. (in prep.; private communication) for a few objects; these do not affect our general conclusions and will be revisited as community lists converge.
@@ 2.3 Gaia+VLBI sample
We introduce a novel methodology that combines high-precision optical astrometry from Gaia with high-resolution VLBI imaging to reliably identify CSO cores. Previous comparisons between Gaia and VLBI astrometry have focused mainly on radio–optical positional offsets in core–jet systems (Petrov & Kovalev, 2017). More recently, exploratory studies have begun to test the joint Gaia+VLBI approach for individual GHz-peaked-spectrum quasars (Kosogorov et al., 2024). Our work constitutes the first systematic application of this technique to a large CSO sample, directly addressing the long-standing difficulty of CSO core identification.
We cross-matched 105 VLBI-selected CSO candidates with Gaia DR3 sources using a conservative search radius and identified 40 robust optical counterparts (Table 4). The Gaia optical centroid probes the immediate vicinity of the accretion disk and the central supermassive black hole (SMBH), thereby providing a stable reference for the AGN core position (see Sect. S2). When registered against VLBI maps with sub-milliarcsecond (mas) resolution, these optical centroids enable precise localization of the central engine, even when the radio core is faint or absorbed. The modest match fraction (38%) further indicates that many CSO nuclei are intrinsically optically faint or heavily obscured.
The cross-matched subset comprises a heterogeneous mix of quasars () and galaxies (). From the redshift distribution (Fig. S1), most CSOs and candidates cluster at low and modrate redshifts (), which is largely set by the Gaia detection limit: galaxies beyond are often too faint for reliable Gaia astrometry. In contrast, core–jet sources show a broader redshift range.
We performed a detailed analysis using the latest DESI Data Release 1 (DESI Collaboration et al., 2025) to assess the detection rates of CSO candidates. Out of our parent sample (105 sources), 63 fall within the DESI footprint, while 42 do not. Among the sources within the DESI footprint, those without Gaia counterparts (35 sources) show a low detection rate () in DESI, with detected sources spanning a redshift range of 0.2–1.5. In contrast, Gaia-detected sources exhibit a higher detection rate () in DESI. This stark difference in detection rates supports the notion that host-galaxy detection is significantly affected by distance. The Gaia non-detections recovered by DESI at higher redshifts confirm that many of these optically faint sources are distant galaxies, not intrinsically faint nearby objects. This result highlights the optical-selection bias and its effect on our CSO sample.
To assess astrometric quality, we screen Gaia solutions using magnitude, the astrometric excess noise (AEN), and its significance (AENS) (Table S2), and we treat significant Gaia–VLBI offsets with caution. Particularly for faint sources ( mag) or extended/structured hosts, low signal-to-noise ratios, asymmetry, or blending can bias the optical centroid and produce spurious Gaia–VLBI offsets, analogous cross-identification uncertainties were reported by Plavin et al. (2019). In such cases, we require independent radio diagnostics before rejecting a CSO classification. This behavior is exemplified by several nearby, well-studied CSOs in Table 1 (e.g., J1407+2827/OQ 208, J1511+0518, J1945+7055), which have strong literature CSO identifications yet show large Gaia–VLBI offsets attributable to Gaia astrometric systematics rather than genuine optical–radio displacements (see also Section 3.1).
This combined approach offers several advantages over traditional, radio-only identification: (i) precise nuclear localization even when the radio core is weak or obscured; (ii) reliable discrimination between CSOs and core–jet sources; (iii) an independent verification of the orientation of the radio structure; and (iv) improved kinematic measurements by anchoring component motions to a fixed, physically motivated reference point.